1/2 t (2/3 \ 3/4)- (4/5 x 5/6)
1
Expert's answer
2013-07-03T08:19:57-0400
1/2 t (2/3 \ 3/4)- (4/5 x 5/6) =1/2 t (2/3 * 4/3)- (4/6)=& 1/2 t (8/9)- (2/3)=1/2 *8/9 t- 2/3=4/9*t-2/3